### v1

@@ 4 Performance Assessment
Scalar i7 gcc 12 libstdc++ 12 Intel MKL 2020
AVX2 i7 clang 15 libstdc++ 12 Intel MKL 2020
AVX512 Xeon gcc 14 libstdc++ 14 Intel MKL 2024
@@
All tests were executed with a single thread for ease of reproducibility. Since it is not possible to control the number of threads used by the Apple Accelerate library on ARM-based Apple platforms, the reported 1-thread speed-ups for high-angular-momentum integrals that use BLAS GEMM may be overly optimistic (such a caveat also applies to the published benchmarks of the SHARK engine22 on Apple platforms).
For a handful of AVX2 tests we also reported double precision performance as percentage of the peak theoretical throughput, the latter obtained by multiplying the clock frequency reported by VTune by 16.
@@
Table 2 reports the performance of our MD implementation for Coulomb (nuclear attraction) integrals due to the potential of 100 point charges. Significant speedups relative to the reference OSHGP implementation in Libint were observed, since even the scalar code is at least 2×2$\times$2 × faster, due to the early summation of point charges possible in the MD framework (Section 2.2.1). The greatest speedups of the scalar code (>13×>13$\times$> 13 ×) are observed for the (s||||s) integrals which only involve the evaluation of Boys functions. Since the interpolation branch of the Boys engine is vectorized even in scalar code (see Section 3.2.1), the additional speedup of SIMD code vs scalar code is due to the benefit of vectorization of the asymptotic branch of the Boys engine. For nonzero $L$ our implementation exhibits substantial (albeit smaller) speedups over the reference OSHGP implementation even using scalar instructions, despite the lower operation count of the latter. For primitive integrals vectorization using AVX2, AVX512, and NEON instruction sets produces speedups over the scalar code of ×$\times 1.6-3.3$ (×$\times 2.4$ average), ×$\times 1.5-4.0$ (×$\times 3$ average), and ×$\times 1.1-1.9$ (×$\times 1.5$ average), respectively. For the contracted integrals vectorization speedups are even higher: $1.7-3.9$ ($2.9$ on average), $1.9-5.7$ ($4$ on average), and $1.1-2.7$ ($1.8$ on average), respectively. Note that the speedups over 2 for contracted (f||||s) integrals with NEON instruction set are somewhat misleading since these figures are relative to the scalar x86-64 code, not to the scalar ARM code. Note that increasing speedup with $L$ is again primarily due to the ability to optimize the early summation (Section 2.2.1) in MD, which allows one to shift more cost outside of the charge summation loop.
@@
Table 3 reports the performance of our MD implementation for 3-center 2-particle integrals. For primitive AOs, the performance of the scalar variant of the new code is substantially superior to that of OS, by ×$\times 2.5$ or more, with few exceptions. Namely, the OSHGPA reference implementation in Libint performs well where ket is large and bra is small, e.g. $(1|66)$, especially if ket is contracted (which should not occur except in atomic natural orbital basis sets where deeply contracted high-$l$ AOs are encountered). SIMD vectorized implementations are always faster than the reference. This is supported by measured FLOP throughputs (averaged over the three cases of contraction depths): evaluation of (p||||pp) and (d||||dd) integrals with LibintX achieved 30% and 50% of peak, respectively, on the AVX2 platform, while the corresponding Libint throughputs were 5% and 7%, respectively. For higher angular momenta even higher throughputs ($\sim 90\%$) are achieved due to the shift of the bulk of the FLOPs to matrix multiplication. The higher FLOP throughput allows our MD-based implementation to overcome the lower operation count of the reference OSHGPA evaluation scheme.
The additional benefit of SIMD vectorization is significant, e.g. on average a factor of 2 speedup is observed with the AVX2 instructions. Note that decreasing SIMD speedup relative to the scalar code is due to the increasing percentage of the cost accounted by the matrix multiplication kernels (which inherently use the SIMD instruction set even in the scalar code). Matrix multiplication also accounts for the anomalously high speedups obtained with the NEON instruction set for high-$L$ integrals (especially, (i||||ii)). The details of Apple Matrix Extensions upon which Apple Accelerate relies are not disclosed at this time. It is not clear if AMX execution units are per core or shared between some cores, or whether such performance is sustained in multithreaded environment. On the other hand, Intel MKL kernels are confined to execute on a single CPU hardware thread.
Note the increased performance when bra is contracted; for example, compare the performance for (p||||ii) integrals with $K_{\text{bra}}=1$ vs $K_{\text{bra}}=5$ with $K_{\text{ket}}=10$. This is due to the early bra contraction optimization (Section 2.2.3).
@@
With increasing angular momentum the relative SIMD performance decreases, due to SIMDized BLAS GEMM accounting for a larger fraction of time, especially where ket is small, as vendor BLAS expectedly beats our bespoke kernel. However, with increasing ket contraction this advantage fades as bra cost decreases in comparison.
@@ 5 Summary
We reported an implementation of the McMurchie-Davidson evaluation scheme for 1- and 2-particle Gaussian AO integrals designed for efficient execution on modern CPUs with SIMD instruction sets. Like in our recent MD implementation for the GPUs1, 3 data parallelism is achieved by computing batches of shellsets of integrals at a time, rather than the one shellset at a time approach popular in the popular CPU-only integral engines like our own Libint engine. By optimizing for the floating point instruction throughput rather than minimizing the number of operations, this approach achieves up to 50% of the theoretical hardware peak FP64 performance for many common SIMD-equipped platforms (AVX2, AVX512, NEON), which translates to speedups of up to $30$ over the state-of-the-art one-shellset-at-a-time implementation of Obara-Saika-type schemes in Libint for a variety of primitive and contracted integrals. As with our previous work, we rely on the standard C++ programming language — such as the std::simd standard library feature to be included in the 2026 ISO C++ standard — without any explicit code generation to keep the code base small and portable. The implementation is part of the open source LibintX library freely available at https://github.com/ValeevGroup/libintx. We invite interested parties to collaborate on the development of fundamental open-source high-performance quantum chemistry building blocks.
@@
Raw benchmark performance data.
@@
This research was supported by the US Department of Energy, Office of Science, via Award DE-SC0019374. The authors acknowledge Advanced Research Computing at Virginia Tech (https://arc.vt.edu/) for providing computational resources and technical support that have contributed to the results reported in this paper.

### v2

@@ 4 Performance Assessment
Scalar i7-9750H (mobile) gcc 12 libstdc++ 12 Intel MKL 2020
AVX2 i7-9750H (mobile) clang 15 libstdc++ 12 Intel MKL 2020
AVX512 Xeon Platinum 8160 gcc 14 libstdc++ 14 Intel MKL 2024
@@
Test Gaussians were generated randomly to mimic typical basis set values and to randomise Boys computations. Libintx timings include time to compute both, integrals and shell pair quantities. Libint timings were obtained by computing a single braket in a loop same number of times as Libintx counterpart with shell pair data initialized outside the loop. In both cases screening was turned off and result buffers were not commited to memory to avoid memory bandwidth artifacts. Tests were repeated 3-4 times with best times reported however the variance was negligible. CPU details are included in supplemental information.
All tests were executed with a single thread for ease of reproducibility. CPU governor was set for performance and frequence scaling was enabled. Since it is not possible to control the number of threads used by the Apple Accelerate library on ARM-based Apple platforms, the reported 1-thread speed-ups for high-angular-momentum integrals that use BLAS GEMM may be overly optimistic (such a caveat also applies to the published benchmarks of the SHARK engine22 on Apple platforms).
For a handful of AVX2 tests we also report double precision performance as percentage of the peak theoretical throughput, the latter obtained by multiplying the clock frequency as reported by VTune by 16 (approximately 64 GFlops peak).
@@
Table 2 reports the performance of our MD implementation for Coulomb (nuclear attraction) integrals due to the potential of 100 point charges. Significant speedups relative to the reference OSHGP implementation in Libint were observed, since even the scalar code is at least $2\times$ faster, due to the early summation of point charges possible in the MD framework (Section 2.2.1). The greatest speedups of the scalar code ($13\times$) are observed for the (s||s) integrals which only involve the evaluation of Boys functions. Since the interpolation branch of the Boys engine is vectorized even in scalar code (see Section 3.2.1), the additional speedup of SIMD code vs scalar code is due to the benefit of vectorization of the asymptotic branch of the Boys engine. For nonzero $L$ our implementation exhibits substantial (albeit smaller) speedups over the reference OSHGP implementation even using scalar instructions, despite the lower operation count of the latter. For primitive integrals vectorization using AVX2, AVX512, and NEON instruction sets produces speedups over the scalar code of $\times 1.6-3.3$ ($\times 2.4$ average), $\times 1.5-4.0$ ($\times 3$ average), and $\times 1.1-1.9$ ($\times 1.5$ average), respectively. For the contracted integrals vectorization speedups are even higher: 1.73.91.7-3.9 ($2.9$ on average), 1.95.71.9-5.7 (44 on average), and 1.12.71.1-2.7 ($1.8$ on average), respectively. Note that the speedups over 2 for contracted (f||s) integrals with NEON instruction set are somewhat misleading since these figures are relative to the scalar x86-64 code, not to the scalar ARM code. Note that increasing speedup with $L$ is again primarily due to the ability to optimize the early summation (Section 2.2.1) in MD, which allows one to shift more cost outside of the charge summation loop.
@@
Table 3 reports the performance of our MD implementation for 3-center 2-particle integrals. For primitive AOs, the performance of the scalar variant of the new code is substantially superior to that of OS, by $\times 2.5$ or more, with few exceptions. Namely, the OSHGPA reference implementation in Libint performs well where ket is large and bra is small, e.g. $(1|66)$, especially if ket is contracted (which should not occur except in atomic natural orbital basis sets where deeply contracted high-$l$ AOs are encountered). SIMD vectorized implementations are always faster than the reference. This is supported by measured FLOP throughputs (averaged over the three cases of contraction depths): evaluation of (p||$p$) and (d||$d$) integrals with LibintX achieved 30% and 50% of peak, respectively, on the AVX2 platform, while the corresponding Libint throughputs were 5% and 7%, respectively. For higher angular momenta even higher throughputs ($\sim 90\%$) are achieved due to the shift of the bulk of the FLOPs to matrix multiplication. The higher FLOP throughput allows our MD-based implementation to overcome the lower operation count of the reference OSHGPA evaluation scheme.
The additional benefit of SIMD vectorization is significant, e.g. on average a factor of 2 speedup is observed with the AVX2 instructions. Note that decreasing SIMD speedup relative to the scalar code is due to the increasing percentage of the cost accounted by the matrix multiplication kernels (which inherently use the SIMD instruction set even in the scalar code). Matrix multiplication also accounts for the anomalously high speedups obtained with the NEON instruction set for high-$L$ integrals (especially, (i||ii)). The details of Apple Matrix Extensions upon which Apple Accelerate relies are not disclosed at this time. It is not clear if AMX execution units are per core or shared between some cores, or whether such performance is sustained in multithreaded environment. On the other hand, Intel MKL kernels are confined to execute on a single CPU hardware thread.
Note the increased performance when bra is contracted; for example, compare the performance for (p||ii) integrals with $K_{\text{bra}}=1$ vs $K_{\text{bra}}=5$ with $K_{\text{ket}}=10$. This is due to the early bra contraction optimization (Section 2.2.3).
@@
With increasing angular momentum the relative SIMD performance decreases, due to BLAS GEMM (internally vectorized) accounting for a larger fraction of time. Since non-SIMD bra transformation (unlike SIMD) can use BLAS GEMM, in cases where bra transform dominates, eg uncontracted $(55|xx)$ and $(66|xx)$ type integrals, non-SIMD code performs comparable or better than SIMD counterpart. However, with increasing ket contraction this advantage fades as bra cost decreases in comparison.
@@
Explicit vectorization does increase the memory required for the intermediates, especially the 2-index Hermite integrals. For example, the $(hh|hh)$ integrals need $286^{2}=81796$ ${\bf\tilde{p}}|{\bf\tilde{q}}$] integrals per vector lane or over 2 MB per AVX2 core, which exceeds the 1MB L3 cache of a typical x86-64 core. This is especially noticable when multiple threads compete for memory bandwidth; this is illustrated in Table 5 where the SIMD $(hh|hh)$ kernel is found to be nearly 2 times slower than the scalar counterpart. Such cases are not typical; e.g., this example assumes that each core executes the same high-$L$ kernel and are unlikely to affect the overall real-world performance; however, if needed these issues are straightforward to resolve.
Performance for the 4-center integrals was also assessed against the established Simint library of Pritchard and Chow10 that emits SIMD instructions by vectorizing over the primitives in the vertical recurrence part of the OSHGP algorithm as well as by batching over multiple shellsets. For all integral classes and contraction degrees the present MD implementation outperforms Simint by a healthy margin. This occurs even for high contraction depth where the advantage of the OSHGP algorithm over the MD algorithm should be most pronounced, and the abundant data parallelism over primitives can be exploited most efficiently by Simint.
@@
$(l_{\mathbf{a}}l_{\mathbf{b}}|l_{\mathbf{c}}l_{\mathbf{d}})$  $\{K_{\text{bra}},K_{\text{ket}}\}=\{1,1\}$ $\{K_{\text{bra}},K_{\text{ket}}\}=\{1,10\}$ $\{K_{\text{bra}},K_{\text{ket}}\}=\{5,10\}$
$(00|00)$ 0.82 0.96 0.93
$(11|11)$ 0.62 0.36 0.34
$(22|22)$ 0.87 0.90 0.85
$(33|33)$ 1.02 1.00 0.93
$(44|44)$ 1.12 0.86 0.72
$(55|55)$ 1.84 1.39 1.27
$(66|66)$ 1.36 1.29 1.20
Table 5: Slowdown of SIMD (I4) MD kernels for 4-center 2-particle integrals relative to the scalar (I1) counterparts while executing on all 6 cores of the Intel i7-9750H CPU. These examples illustrate the worst-case performance degradation due to the SIMD-driven memory pressure increase.
$(l_{\mathbf{a}}l_{\mathbf{b}}|l_{\mathbf{c}}l_{\mathbf{d}})$  $\{K_{\text{bra}},K_{\text{ket}}\}=\{1,1\}$ $\{K_{\text{bra}},K_{\text{ket}}\}=\{1,10\}$ $\{K_{\text{bra}},K_{\text{ket}}\}=\{5,10\}$
$(00|00)$ 1.78 1.71 1.52
$(10|00)$ 1.92 2.20 1.86
$(11|00)$ 1.64 2.41 2.01
$(11|11)$ 1.24 2.02 1.96
$(22|00)$ 1.15 1.83 1.77
$(22|11)$ 1.22 1.99 1.86
$(22|22)$ 1.95 2.27 2.07
$(33|00)$ 1.75 1.96 1.52
$(33|11)$ 2.02 3.71 3.19
$(33|33)$ 3.33 2.72 2.50
$(44|00)$ 1.60 2.18 1.65
$(44|11)$ 2.22 5.02 4.43
$(44|44)$ 3.10 2.22 1.90
$(55|00)$ 4.67 4.44 1.91
$(55|11)$ 2.32 4.52 3.31
$(55|55)$ 4.96 2.30 1.65
Table 6: Speedup of the MD implementation of Coulomb 4-center 2-particle integrals in LibintX vs the SIMD-optimized OSHGP implementation in Simint (Ref. 10).a
a Performance was measured on a single AVX2 core of the Intel i7-9750H CPU.
@@ 5 Summary
We reported an implementation of the McMurchie-Davidson evaluation scheme for 1- and 2-particle Gaussian AO integrals designed for efficient execution on modern CPUs with SIMD instruction sets. Like in our recent MD implementation for the GPUs1, 3 data parallelism is achieved by computing batches of shellsets of integrals at a time, rather than the one shellset at a time approach popular in the popular CPU-only integral engines like our own Libint engine.
By optimizing for the floating point instruction throughput rather than minimizing the number of operations, this approach achieves up to 50% of the theoretical hardware peak FP64 performance for many common SIMD-equipped platforms (AVX2, AVX512, NEON), which translates to speedups of up to $30$ over the state-of-the-art one-shellset-at-a-time implementation of Obara-Saika-type schemes in Libint for a variety of primitive and contracted integrals. In this context the relative GPU speedups are much more modest, e.g., 28, 46, and 83 rather than 592, 223, and 1172 for uncontracted $(11|00)$, $(22|22)$, $(66|66)$, respectively, when comparing V100 numbers from our prior work and i7-9750h numbers here.
As with our previous work, we rely on the standard C++ programming language — such as the std::simd standard library feature to be included in the 2026 ISO C++ standard — without any explicit code generation to keep the code base small and portable. The implementation is part of the open source LibintX library freely available at https://github.com/ValeevGroup/libintx. We invite interested parties to collaborate on the development of fundamental open-source high-performance quantum chemistry building blocks.
@@
Raw benchmark performance data, platform information.
@@
This research was supported by the US Department of Energy, Office of Science, via award DE-SC0022327. The authors acknowledge Advanced Research Computing at Virginia Tech (https://arc.vt.edu/) for providing computational resources and technical support that have contributed to the results reported in this paper.